### v2

# Understanding and Bridging the Planner-Coder Gap: A Systematic Study on the Robustness of Multi-Agent Systems for Code Generation

Abstract.
Multi-agent systems (MASs) have emerged as a promising paradigm for automated code generation, demonstrating impressive performance on established benchmarks by decomposing complex coding tasks across specialized agents with different roles. Despite their prosperous development and adoption, the fundamental mechanisms underlying their robustness remain poorly understood, raising critical concerns for real-world deployment. This paper conducts a systematic empirical study to uncover the internal robustness flaws of MASs using an automated mutation-based methodology. By designing a testing pipeline incorporating semantic-preserving mutation operators and a novel fitness function, we assess mainstream MASs across multiple datasets and LLMs. Our findings reveal substantial robustness flaws: semantically equivalent inputs cause drastic performance drops, with MASs failing to solve 7.9%–83.3% of problems they initially resolved successfully after applying the semantic-preserving mutations.
Through comprehensive failure analysis, we discover a fundamental cause underlying these robustness issues: the planner-coder gap, which accounts for 75.3% of failures. This gap arises from information loss in the multi-stage transformation process where planning agents decompose requirements into underspecified plans, and coding agents subsequently misinterpret intricate logic during code generation. Based on this formulated information transformation process, we propose a repairing method that mitigates information loss through multi-prompt generation and introduces a monitor agent to bridge the planner-coder gap. Evaluation shows that our repairing method effectively enhances the robustness of MASs by solving 40.0%–88.9% of identified failures. Re-execution of the testing process on the repaired MASs shows that the number of detected failures decreases up to 85.7%, demonstrating that repaired MASs exhibit superior robustness. Our work uncovers critical robustness flaws in MASs and provides effective mitigation strategies, contributing essential insights for developing more reliable MASs for code generation.
multi-agent systems, code generation, robustness
conference: The ACM International Conference on Software Testing and Analysis; 3-9 Oct, 2026; Oakland, California, United Statesccs: Software and its engineering Software testing and repairingccs: Software and its engineering Robustness
1. Introduction
The advent of large language models (LLMs) has fundamentally transformed code generation methodologies, enabling the automated production of code that fulfills user requirements expressed in natural language (Jiang et al., 2024a; Hou et al., 2024). Among these methodologies, multi-agent systems (MASs) have emerged as a particularly promising paradigm. By leveraging the same backend LLM, MASs decompose complex coding tasks into manageable sub-tasks and distribute them among specialized agents with distinct roles (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a; Islam et al., 2024, 2025). For instance, planning agents design the overall coding logic and algorithmic structure, while coding agents implement the actual code based on these plans. Through carefully crafted prompts and communication protocols, these specialized agents collaborate effectively, achieving impressive performance across diverse coding challenges on established benchmark datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a).
Despite their demonstrated effectiveness on benchmarks, a critical limitation exists in our understanding of MAS robustness for code generation. While MASs work well on standard evaluation based on test cases, the fundamental mechanisms underlying their failures remain poorly understood—particularly when faced with semantically equivalent variations of input requirements. This “black box” problem poses significant risks for real-world deployment: non-robust MASs may generate inconsistent or incorrect code solutions, resulting in system failures and increased debugging costs (Mastropaolo et al., 2023; Zhong et al., 2024). Without a systematic understanding of why and how MASs fail under semantic-preserving input variations, it remains difficult to develop principled solutions for enhancing their reliability.
To address this knowledge gap, we conduct a systematic empirical study to uncover the internal robustness flaws of MASs using an automated mutation-based methodology. Inspired by existing efforts on AI model robustness testing (Wang et al., 2023a; Mastropaolo et al., 2023; Chen et al., 2021b; Iakusheva et al., 2023; Xie et al., 2022; Rehman and Srinivasan, 2023), we design four semantic-preserving mutation operators that systematically transform input questions while preserving their intended semantics. Our testing methodology is guided by a novel fitness function that evaluates both the quality of generated code and the semantic similarity of generated plans, providing fine-grained feedback for mutation selection. We assess three mainstream MASs (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a) across multiple datasets and backend LLMs. Our empirical findings reveal substantial robustness flaws: semantically equivalent inputs cause drastic performance drops, with MASs failing to solve 7.9%–83.3% of problems they initially resolved successfully after applying semantic-preserving mutations.
Through comprehensive failure analysis of the testing results, we discover a fundamental cause underlying these robustness issues: the planner-coder gap, which accounts for 75.3% of all observed failures. This gap arises from information loss in the multi-stage transformation process inherent to MAS architectures. Specifically, planning agents decompose user requirements into high-level plans that are often logically sound but lack sufficient implementation details, while coding agents subsequently misinterpret intricate logic or ambiguous expressions during code generation. We systematically characterize this gap through five distinct error patterns (EPs): EP-1: Gap in Core Concepts, where coding agents misunderstand fundamental algorithmic concepts in plans; EP-2: Gap in Edge Cases, where boundary conditions specified in plans are overlooked during implementation; EP-3: Gap in Complex Logic, where multi-step reasoning or nested conditions are incorrectly translated; EP-4: Gap in Relational Phrases, where quantity or spatial relationships are misinterpreted; and EP-5: Gap in Condition Judgments, where conditional branches or guard clauses are improperly implemented. These patterns reveal that enhancing MAS robustness fundamentally requires bridging the planner-coder gap to mitigate information loss across agents.
Based on our formulated understanding of the information transformation process, we propose a repairing method to mitigate information loss and bridge the planner-coder gap. Our approach incorporates two synergistic components. First, multi-prompt generation diversifies input question expressions through semantic-preserving mutations, reducing the likelihood of misinterpretation. Second, we introduce a specialized monitor agent as an intermediary between planning and coding agents. This monitor agent operates through two key processes: (1) a plan interpretation process that provides detailed explanations to enhance the coding agent’s comprehension, and (2) a code validation process that verifies alignment between generated code and the interpreted plan. By explicitly addressing information loss at the critical planner-coder interface, our monitor agent effectively bridges the communication gap underlying most robustness failures.
We conduct extensive experiments to evaluate our repairing method across three prominent MASs (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a) with three backend LLMs (Achiam et al., 2023; Zhu et al., 2024; OpenAI, 2022) on four datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a). Results demonstrate that our approach successfully repairs 40.0%–88.9% of identified failures. When we re-execute the testing process on repaired MASs, discovered failures decrease by up to 85.7%, providing empirical evidence of substantially improved robustness. Ablation studies confirm that both components are necessary, as they address complementary aspects of the robustness challenge. Overall, we summarize our contributions as follows:
  • We conduct the first systematic empirical study on the robustness of MASs for code generation, focusing on their behavior under semantically equivalent input variations. Our automated mutation-based methodology incorporates semantic-preserving mutations and a novel fitness function to effectively assess mainstream MASs, revealing substantial robustness flaws where performance drops by 7.9%–83.3% under semantic-preserving transformations.
  • We discover and characterize the planner-coder gap as a fundamental cause of MAS robustness failures, accounting for 75.3% of observed issues. Through systematic analysis, we identify five distinct error patterns that manifest this gap and formulate the multi-stage information transformation process in MAS code generation, providing theoretical grounding for understanding information loss across agent boundaries.
  • Based on our empirical findings and theoretical formulation, we propose a repairing method that mitigates information loss through multi-prompt generation and monitor agent insertion. Comprehensive experiments demonstrate that our approach repairs 40.0%–88.9% of identified failures and reduces failure rates by up to 85.7% in subsequent testing, validating its effectiveness in enhancing MAS robustness.
2. Background
Code generation has been a central focus of software engineering research, aiming to automate programming tasks and enhance developer productivity by generating desired code from natural language specifications (Shin and Nam, 2021; Dehaerne et al., 2022; Cao et al., 2024; Du et al., 2024; Mathews and Nagappan, 2024; Gao et al., 2024). MASs are artificial intelligence systems that leverage LLMs as their core computational engines to simulate domain experts with specialized capabilities (Qian et al., 2024; Hong et al., 2024; Zhang et al., 2024a). With advances in prompt engineering (Liu et al., 2023), developing MASs to enhance code generation has emerged as a prominent research direction (He et al., 2024; Tran et al., 2025).
A prevalent multi-agent approach to code generation employs role specialization and iterative feedback mechanisms to provide code snippets that satisfy the user input (Liu et al., 2024). As shown in Fig. 1, the typical architecture for code generation MASs consists of three primary parts: planning, coding, and feedback-driven refinement, which are typically implemented through planner, coder, and tester (He et al., 2024).
Figure 1. Pipeline of MASs for code generation.
Fig. 1 depicts a typical workflow of a code generation MAS. Upon receiving a natural language requirement, the system routes the input to the planning agent (planner) to generate a clear and structured coding plan. Note that the planning agent does not simply paraphrase the user input. Instead, it serves as an important coding step that decomposes the coding problem into a set of requirements and designs several key logical steps to solve the problem, which determines the algorithm and the logic for code implementation (Zhang et al., 2024a; Dong et al., 2024). The coding agent (coder) then implements the solution following the provided plan. Subsequently, the testing agent (tester) automatically generates test cases and evaluates the implementation, producing a test report. If the code fails, the report is returned to the coder for iterative refinement until the solution passes all tests or the computational budget is exhausted.
This specialized multi-agent workflow represents a fundamental paradigm shift from single-LLM code generation approaches. Unlike the end-to-end paradigm where a unified model directly transforms natural language into executable code (Wang et al., 2021; Fried et al., 2023), MASs decompose the complex task into specialized subtasks handled by domain-specific agents. This paradigm demonstrates remarkable advantages in handling complex coding tasks. Nevertheless, our analysis in Sec. 5 reveals critical robustness issues concerning the interaction between planner and coder.
3. Motivation
MASs (Hong et al., 2024; Dong et al., 2024; Zhang et al., 2024a) have emerged as a transformative technology with immense commercial potential. IT companies have invested significant effort in the development of MASs, with Google Cloud announcing comprehensive enhancements to Vertex AI that include the Agent Development Kit (ADK) and Agent Engine for building and deploying enterprise-grade MASs at scale (50). OpenAI has also launched new APIs and tools specifically designed for agentic applications, including the Responses API and the Agents SDK for orchestrating multi-agent workflows, demonstrating growing commercial interest in MASs (38).
However, assessing the reliability and robustness of these complex systems remains a significant research gap. Unlike single agents and individual models, MASs involve intricate inter-agent communications, dynamic role assignments, and collaborative decision-making processes that make their behavior difficult to predict and evaluate systematically (He et al., 2024). This gap in assessment capabilities becomes particularly concerning given the high-stakes commercial applications these systems are designed to serve (He et al., 2024). Inspired by previous work that employs testing to assess the robustness of AI systems (Li et al., 2023; Mastropaolo et al., 2023), we assess the robustness of MASs for code generation based on semantically equivalent input mutations. In particular, we presume that a robust MAS should maintain consistency when addressing semantically equivalent input and generating correct programs regardless of how developers articulate their requirements. This idea stems from the observation that developers’ requests typically vary in vocabulary and phrasing due to the variability of natural language.
To illustrate this idea, consider a simple example where a developer might request: (1) “return the n-th Fibonacci number”, (2) “give back the n-th number in the Fibonacci sequence.” Although these requests use different phrasing, their semantic meaning is identical. A robust MAS should generate functionally correct solutions for both expressions. However, we find that even the state-of-the-art MASs suffer from inconsistencies in handling such semantically equivalent expressions. For example, PairCoder (Zhang et al., 2024a) generates correct code for the first request, while failing on the second due to incorrect edge case handling.
4. Fuzzing MAS with Semantically Equivalent Requirements
To fill the gap in understanding of MASs’ robustness in code generation as discussed in Sec. 3, we design a fuzz testing framework specifically adapted to MAS architectures. Our approach systematically generates semantically equivalent variations of programming requirements to expose robustness flaws arising from inter-agent communication—a vulnerability dimension not addressed by traditional testing approaches. A key innovation is our dual-component fitness function, which evaluates mutations along two dimensions: (1) final code quality degradation and (2) semantic drift in intermediate plans. This MAS-specific design enables us to systematically uncover the planner-coder gap, a fundamental structural flaw that represents the dominant failure mode in contemporary MASs (see Sec. 5).
4.1. Fuzzing Methodology
We first present four semantic-preserving mutation operators that generate alternative expressions of selected questions. Then, we introduce a novel fitness function that calculates rewards based on the differences in both output code and generated plans between the original and mutated questions, to provide comprehensive guidance for subsequent fuzzing iterations.
Mutation Operators. Our mutation approach focuses on generating diverse expressions while preserving the semantic meaning of input questions. To achieve this, we design four sentence-level mutation operators that modify the expression styles of natural language requirements without altering the original semantic content.
Table 1. Semantic-preserving mutation operators and descriptions.
Operator Description Rephrase Instruct an LLM to rewrite a sentence using other words while maintaining the overall meaning. Insert Prompt an LLM to append one additional sentence at the end of the description based on the semantic content. Expand Use an LLM to expand one sentence into two by distributing its semantic content. Condense Apply an LLM to condense two consecutive sentences into one sentence using appropriate conjunctions.
Table 1 presents these four mutation operators: Rephrase, Insert, Expand, and Condense. We implement these operators using GPT-4o (Achiam et al., 2023), a state-of-the-art LLM, through carefully designed prompts. This design effectively addresses two key challenges from previous work (Wang et al., 2023a; Mastropaolo et al., 2023): (1) it minimizes semantic deviations that may occur during rule-based mutation, and (2) it produces natural and fluent expressions in the mutated questions. Moreover, as shown in prior work (Yu et al., 2024b), LLM-generated mutations effectively simulate diverse human expressions. Our operators cover prevalent semantic-preserving transformations (Yu et al., 2024b), providing comprehensive coverage of semantic variations.
To confirm the reliability of our mutation operators in terms of semantic preservation, we conduct manual verification on 500 randomly sampled mutants, confirming that 99.2% of the mutants preserve semantics. Additionally, we measure the semantic similarity of all mutants to the original questions using Sentence-Transformers (Reimers and Gurevych, 2019), whose results show 98.4% average similarity.
Fitness Function. To determine whether a mutated question should be retained in the seed pool for further mutation, we propose a novel fitness function that calculates rewards for MAS outputs on mutated questions. Unlike previous coverage-driven fuzzing approaches (Zalewski, 2025; Böhme et al., 2016; Ognawala et al., 2018; Veggalam et al., 2016) on software, our fitness function considers the multi-agent nature of MASs by comparing the output quality between the seed questions and their mutated variants across two distinct aspects: code and plan.
Code Reward. To mitigate the occasional failure due to randomness rather than underlying robustness issues, we execute MASs for $n$ times for each question. We collect the generated code and calculate the rate of completions passing all test cases provided in the dataset. Based on the insight that an increasing number of failure trials on the mutated questions indicate degraded MAS performance (Wang et al., 2023b; Chen et al., 2021a), we calculate the code reward by measuring the difference between the pass rates of code generated from original versus mutated questions:
(1) $\mathcal{R}_{C}=\frac{1}{n}\left(\sum_{i=1}^{n}c_{i}-\sum_{i=1}^{n}\hat{c}_{i}\right)$
where $n$ represents the number of attempts, $c_{i}$ and $\hat{c}_{i}$ indicate whether the $i$-th generation passes the test (1 for pass, 0 for fail). Positive rewards indicate that the mutated question reduces MAS performance, making it valuable for identifying robustness issues in subsequent fuzzing processes.
Plan Reward. Apart from the generated code, we also compare the semantic of plans generated by the planning agent for the original and mutated questions. Significant changes in the generated plans often indicate potential robustness issues in the planning phase, which substantially influence the coding agent’s ability and the final code quality. We apply Sentence-Transformers (Reimers and Gurevych, 2019) to obtain the embedding and calculate semantic similarity between corresponding plans. The plan reward is computed by subtracting the normalized similarity from 1:
[…]
where $p_{i}$ and $\hat{p}_{i}$ represent the vector embeddings of plans generated from the seed question and the mutated question, respectively, for the $i$-th generation.
The final fitness function is calculated as the sum of the code reward and the plan reward. This combination enables measuring robustness issues in both the code quality degradation and semantic drift in the planning stage, which is the key to comprehensively understanding MAS robustness.
Input: Multi-Agent System $M$, Seed Pool $\mathcal{S}$, Fitness Function $\mathcal{F}$, Budget $\mathcal{B}$, Attempts $n$
1
Output: Failure Question Set $\mathcal{R}$
2
[…]
4
5while $ConsumedQueries<\mathcal{B}$ and $\mathcal{S}\neq\emptyset$ do
6  $seed\leftarrow\text{MCTS-exploring}(\mathcal{S})$ ;
  $mutated\_seed\leftarrow\text{Mutate}(seed);$ // Mutate the seed question with randomly selected operator
7
8 $failures\leftarrow 0$;
9  for $i=1$ to $n$ do
     $result\leftarrow\text{Execute}(M,mutated\_seed);$ // Execute the MAS with the mutated seed
10     if $result=\text{False}$ then
11        $failures\leftarrow failures+1$;
12
13
14
15 if $failures=n$ then
     $\mathcal{R}\leftarrow\mathcal{R}\cup\{mutated\_seed\};$ // Add the mutated seed to the failure set
     $original\_seed\leftarrow\text{GetParent}(seed);$ // Get the original seed question
     $\mathcal{S}\leftarrow\mathcal{S}\setminus\{original\_seed\};$ // Stop fuzzing on this branch
16
17 else
     $reward\leftarrow\mathcal{F}(result);$ // Calculate reward using fitness function
18     if $reward>0$ then
        $\mathcal{S}\leftarrow\mathcal{S}\cup\{mutated\_seed\};$ // Add the mutated seed to the seed pool
19
20
21
return $\mathcal{R}$
Algorithm 1 Fuzzing Pipeline
Fuzzing Pipeline. Following prior fuzzing methodologies (Fioraldi et al., 2022; Zalewski, 2025), our approach encompasses four major components: seed pool initialization, seed selection, mutation, and execution. The workflow of our fuzzing process is presented in Alg. 1. The algorithm takes the target MAS $\mathcal{M}$, seed pool $\mathcal{S},$ fitness function $\mathcal{F}$, fuzzing budget $\mathcal{B}$ and total attempts $n$ as inputs, and outputs failure question set $\mathcal{R}$. Specifically, we first create an initial seed pool from questions in established datasets that the target MAS successfully resolves. For each fuzzing iteration, following recent advances in fuzzing methodologies, we employ MCTS-exploring (Yu et al., 2024b) as our seed selection mechanism (line 3), which effectively identifies promising seeds while avoiding local optima. Once a seed is selected, we randomly choose one of our four semantic-preserving mutation operators to generate a semantically equivalent version of the question (line 4). Subsequently, the mutated question is submitted to the MAS, which executes the code generation process $n$ times (where $n=10$ in our implementation) and records the failure numbers (lines 6–9). If the MAS fails to correctly solve the mutated question in all trials, we identify this as an unsolvable variant and halt further fuzzing on that branch (lines 11–13). If the MAS successfully handles the mutated question, we calculate the reward using our fitness function to determine whether to retain the mutation for further exploration (lines 15–17). We continue the above process until we reach predefined resource constraints or exhaust valid seeds for mutation.
4.2. Experiment Setup
MASs. We evaluate three widely-adopted code generation MASs: Self-Collaboration Code Generation (Dong et al., 2024) (abbr. SCCG), MetaGPT (Hong et al., 2024), and PairCoder (Zhang et al., 2024a). All systems incorporate agents responsible for planning and coding. MetaGPT represents one of the earliest MAS frameworks, applying five agents to collaboratively develop software based on user instructions. SCCG proposes the workflow of planning, coding and testing, building a popular framework for many subsequent works (Zhang et al., 2024a; Islam et al., 2024). PairCoder is the state-of-the-art MAS that includes plan selection.
Backend LLMs. MASs employ the same backend LLM for all constituent agents. We use three different LLMs as backend models for the aforementioned MASs: GPT-3.5-Turbo (OpenAI, 2022) (abbr. GPT-3.5), GPT-4o (Achiam et al., 2023), and Deepseek-Coder-V2 (Zhu et al., 2024) (abbr. Deepseek). All these LLMs are widely adopted and demonstrate strong capabilities in code generation (Joel et al., 2024; Fakhoury et al., 2024; Jiang et al., 2024a). We apply the same parameters for the backend LLMs as the original MAS papers (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a).
Datasets. We employ four code generation datasets: HumanEval ET (Chen et al., 2021a) (abbr. HumanEval), MBPP ET (Austin et al., 2021) (abbr. MBPP), CodeContest (Li et al., 2022) and CoderEval (Yu et al., 2024a), which are widely used for benchmarking MASs for code generation (Dong et al., 2024; Lin et al., 2025). HumanEval and MBPP (Dong et al., 2025) provide well-constructed coding problems, serving as standard benchmarks. CodeContest (Li et al., 2022) represents competitive programming scenarios, offering more challenging problems. Notably, CoderEval (Yu et al., 2024a) is latest dataset that mirrors real-world development with data sourced from actual GitHub repositories, thereby offering a modern and realistic assessment of MASs. These datasets, ranging from classical benchmarks to cutting-edge real-world scenarios, are chosen for evaluation across various coding contexts.
Dataset Splitting. It is important to note that we apply 50% of the data in each dataset for testing, and the rest for repairing (see Sec. 6). In other words, this strictly prohibits using the same data for both fuzzing and repairing in order to avoid data leakage. We use the sanitized version of MBPP, the test set of CodeContest, and the standalone-level of CoderEval to save resources.
Evaluation Metric and Fuzzing Parameters. We adopt Pass@k (Chen et al., 2021a) as our primary metric to comprehensively assess MAS robustness. Following previous works (Chen et al., 2021a; Li et al., 2022), we set $k=10$ in our evaluation. Compared to Pass@1, Pass@10 mitigates the influence of occasional failures due to randomness and can reflect more nuanced performance differences among MASs (Austin et al., 2021). To balance comprehensiveness and resource consumption, we set the query budget of our fuzzing to 10,000. For the MCTS-Explore algorithm, we apply identical parameter settings as specified in (Yu et al., 2024b). To prevent local optima, after a question is selected 15 times, we terminate fuzzing on the branch.
4.3. Fuzzing Results
Table 2. Performance comparison on the original performance and performance after fuzzing.
MAS Backend LLM HumanEval MBPP CodeContest CoderEval
Fuzzing Original Drop Fuzzing Original Drop Fuzzing Original Drop Fuzzing Original Drop
SCCG GPT-3.5 0.4815 0.7160 $\downarrow$ 32.6% 0.4590 0.5995 $\downarrow$ 23.4% 0.0364 0.1091 $\downarrow$ 66.7% 0.2540 0.4286 $\downarrow$ 40.7%
GPT-4o 0.6975 0.8395 $\downarrow$ 16.9% 0.5761 0.6909 $\downarrow$ 16.6% 0.2121 0.3394 $\downarrow$ 37.5% 0.3016 0.4921 $\downarrow$ 38.7%
Deepseek 0.5617 0.7345 $\downarrow$ 23.5% 0.4964 0.6440 $\downarrow$ 22.9% 0.0242 0.0970 $\downarrow$ 75.0% 0.1429 0.3810 $\downarrow$ 62.5%
MetaGPT GPT-3.5 0.4085 0.7073 $\downarrow$ 42.2% 0.5761 0.6651 $\downarrow$ 13.4% 0.0242 0.0788 $\downarrow$ 69.2% 0.3016 0.4127 $\downarrow$ 26.9%
GPT-4o 0.6707 0.8170 $\downarrow$ 17.9% 0.6089 0.6979 $\downarrow$ 12.8% 0.1879 0.3091 $\downarrow$ 39.2% 0.3333 0.4603 $\downarrow$ 27.6%
Deepseek 0.5915 0.7378 $\downarrow$ 19.8% 0.5972 0.6812 $\downarrow$ 12.9% 0.0242 0.1515 $\downarrow$ 83.3% 0.2540 0.3968 $\downarrow$ 36.0%
PairCoder GPT-3.5 0.6829 0.7744 $\downarrow$ 11.8% 0.6721 0.7307 $\downarrow$ 8.0% 0.1030 0.2364 $\downarrow$ 43.6% 0.0952 0.4127 $\downarrow$ 76.0%
GPT-4o 0.7622 0.8475 $\downarrow$ 10.1% 0.6791 0.7377 $\downarrow$ 7.9% 0.4424 0.5273 $\downarrow$ 16.1% 0.1587 0.4762 $\downarrow$ 66.7%
Deepseek 0.7683 0.8536 $\downarrow$ 10.0% 0.6417 0.7025 $\downarrow$ 8.7% 0.1576 0.2788 $\downarrow$ 43.5% 0.2698 0.4603 $\downarrow$ 41.4%
Table 2 presents the performance differences of MASs before and after fuzzing. The Drop represents the decrease rate of pass@10 after fuzzing. Our results reveal that all MASs with different backend LLMs experience performance degradation, with pass@10 decreasing by 7.9%– 83.3%. Specifically, SCCG (GPT-3.5) on CodeContest demonstrates the highest accuracy drop, while PairCoder (GPT-4o) on MBPP exhibits the best robustness. These findings demonstrate that current MASs suffer from robustness issues when confronted with semantically equivalent questions, despite achieving promising results on original datasets.
Our fuzzing results reveal that different MASs exhibit varying degrees of robustness issues. PairCoder consistently demonstrates the best robustness across most datasets, while SCCG and MetaGPT suffer from more severe performance degradation. For instance, when using GPT-3.5 as the backend LLM on HumanEval , PairCoder’s performance drops by only 11.8%, whereas SCCG and MetaGPT experience over 30% decline. The superior robustness of PairCoder can be attributed to its navigator agent that generates and selects optimal plans from multiple candidates. However, PairCoder’s advantage diminishes on CoderEval, where it suffers the largest performance drop due to its heavy reliance on public test cases for iterative code refinement, which CoderEval does not provide. Meanwhile, MetaGPT shows better robustness on MBPP and CoderEval, which contain shorter questions with minimal explanations, aligning with MetaGPT’s design philosophy of processing brief user requirements. These findings suggest that plan selection mechanisms and dataset characteristics may influence MAS robustness.
Moreover, we notice that MASs exhibit more severe robustness issues when dealing with longer and more complex questions. Among the four datasets, CodeContest and CoderEval exhibit substantially larger performance drops than HumanEval and MBPP. This pattern can be explained by their distinct characteristics: CodeContest contains competition-level problems with lengthy descriptions and strict input format constraints, while CoderEval involves intricate problem settings that mirror real-world scenarios. Our findings reveal that MASs are particularly sensitive to these questions, where even minor perturbations can trigger failures. This heightened sensitivity raises concerns about the reliability of MASs in practical applications.
We further observe that the capability of the backend LLM affects MAS robustness, with powerful LLMs demonstrating superior robustness. Across all datasets, MASs powered by GPT-3.5 experience more significant performance degradation than those using GPT-4o. We attribute this difference to the limited semantic understanding of weaker LLMs, which makes them less capable of accurately interpreting diverse expressions and thus more vulnerable to robustness issues.
Our Findings: Existing MASs generally suffer from robustness issues, manifesting a substantial performance drop when processing semantically equivalent questions. This problem is further exacerbated by longer and more complex questions, but can be mitigated by using more powerful backend LLMs. Furthermore, our results reveal that plan selection and refinement mechanisms may enhance MAS robustness, indicating a potential direction to improve MAS design.
5. Implications from Failure Cases
The substantial robustness flaws revealed in Sec. 4—where semantically equivalent mutations cause performance drops of 7.9%–83.3%—raise a fundamental question: what underlying mechanisms cause MASs to fail on inputs they should handle correctly? To answer this question, we conduct a systematic failure analysis on over 700 uncovered failures. Specifically, we randomly sample 20% of failures and manually examine both the generated code and internal outputs (e.g., plans) to investigate the root causes. To mitigate subjectivity, two software developers first build a failure taxonomy through a pilot study on 5% of cases, then independently categorize the remaining cases. The final Cohen’s Kappa score (Cohen, 1960) is 0.88, indicating high agreement.
Table 3. Distribution of reasons for failures.
Category #Failure Percentage
Planner-Coder Gap 113 75.3%
Plan Logic Errors 23 15.3%
Invalid 14 9.3%
Our systematic analysis yields a critical insight: the dramatic performance drops observed in Sec. 4.3 (7.9%–83.3%) are not random occurrences, but stem from a fundamental structural flaw in MAS architecture—the planner-coder gap. As presented in Table 3, 75.3% of failures are rooted in this gap, where semantic alignment breaks down during multi-stage information transformation from planning to coding agents. Specifically, generated plans maintain logical correctness, but coding agents fail to properly implement correct code from these plans. This gap—characterized by information loss and semantic drift across agent boundaries—explains why semantically equivalent mutations, which should not affect correctness, instead trigger systematic failures. In contrast, only 24.7% of failures stem from other causes: 15.3% from plan logic errors (atomic LLM errors where planners generate algorithmically incorrect plans), and 9.3% from invalid cases (inherent ambiguities in original requirements). The planner-coder gap thus represents the dominant, addressable failure mode in contemporary MASs.
This finding uncovers a previously overlooked dimension that extends beyond the current research focus. While preliminary studies (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2024) have established that planning constitutes a critical coding step determining implementation logic and that effective inter-agent communication is essential for correct plan implementation, most have concentrated primarily on generating algorithmically correct plans. Our analysis, however, demonstrates that even with logically correct plans, the planner-coder gap emerges as a fundamental impediment to robust code generation in MASs. This gap manifests as a breakdown in semantic alignment during the multi-stage information transformation process, where the planner’s high-level abstractions fail to preserve sufficient implementation constraints for the coder to reconstruct the original intent.
Through comprehensive analysis, we identify two principal mechanisms underlying this gap: (1) semantic drift, where the planner’s intent is progressively diluted through abstraction, resulting in plans that are logically sound but lack critical implementation details; and (2) context fragmentation, where the coder loses access to crucial constraints and boundary conditions that were implicit in the original requirements but not explicitly preserved in the plan. Consequently, we systematically categorize the planner-coder gap into five distinct error patterns (EPs) and analyze their respective distribution frequencies.
EP-1. Gap in Core Concepts (32.7%).:
Many failure questions involve specific core concepts essential to problem-solving. For example, the requirement of HumanEval-26111Short for “the 26th question in HumanEval.” Similar notations are used in the rest of the paper. is to “remove all duplicates in the input list”. Here, “remove duplicates” means deleting all elements that appear more than once. The planner provides correct instructions to “remove all duplicates” but does not clarify the meaning of “duplicate”. This leaves a gap that the coder misinterprets “remove duplicate” as removing repeated occurrences of elements, but keeping one instance of all duplicates.
EP-2. Gap in Edge Cases (19.5%).:
In some cases, the generated code returns incorrect results for edge cases. Although the plan correctly identifies edge cases using expressions such as “handle the cases where…”, the coder still fails to handle them when there are no clear explanations of the expected results for edge cases, revealing the gap in edge case handling.
EP-3. Gap in Complex Logic (15.9%).:
Coders can fail to follow the logic when there are insufficient concrete explanations and analysis. In some cases, steps in the plan contain complex logic like “sort the coordinates of each row by columns”, which the coder fails to comprehend.
EP-4. Gap in Relational Phrases (9.7%).:
Our case study demonstrates that coders are prone to misunderstand phrases that express quantity or spatial relationships between variables, indicating the need for further interpretation of these expressions. Expressions like “at least as much as”, “repeated two multiply two times” are often misunderstood by the coder in our experiments.
EP-5. Gap in Condition Judgments (22.1%).:
Many questions require different logic paths for various situations. Code implementations sometimes omit certain logic paths specified in the plan when there are no detailed explanations of the condition for each logic path. For example, in HumanEval-128, the plan for the mutated question correctly illustrates the proper logic as “calculate the product of signs and sum of all numbers”. However, the question requires special treatment when the number is zero. Without a detailed explanation for this requirement in the plan, the coder omits the logic path to check if the number is zero.
Implications for Robustness Enhancement. Each of these error patterns highlights the structural nature of the planner-coder gap: the patterns are not attributable to prompt understanding failures at the planning stage, but rather stem from communication breakdowns between agents where the planner’s high-level abstractions fail to preserve sufficient implementation details. Accounting for 75.3% of observed failures, the planner-coder gap represents the dominant robustness issue in MASs. This finding motivates our repair approach in Sec. 6, which targets this gap through mechanisms that bridge the planner-coder interface.
6. Augmenting MAS Robustness
Having identified the planner-coder gap as the root cause of 75.3% of robustness failures, we now address the question: how can we mitigate this fundamental structural flaw? To answer this, we first formulate the planner-coder gap by modeling the multi-stage information transformation process in MASs. This formulation reveals that information loss accumulates across agents, providing a theoretical foundation for improvement. Based on this formulated understanding, we propose a repairing method comprising two components: multi-prompt generation and monitor agent insertion.
6.1. Reflection and Formulation of Planner-Coder Gap
Our empirical discovery in Sec. 5 reveals that the planner-coder gap is not merely a prompt engineering issue, but a fundamental vulnerability in the multi-stage transformation architecture of MASs. To design an effective repair, we must first understand why this gap occurs. Unlike single LLMs that directly encode input descriptions into code, MASs employ a multi-agent workflow where the planner plays a central role in determining generation logic (Zhang et al., 2024a; Hong et al., 2024; Dong et al., 2024; Islam et al., 2024). We formulate this process to characterize where and how information loss occurs.
In MASs, the planner serves as the central orchestrator, deciding the logic of the code to generate (Jiang et al., 2024b; Zhang et al., 2024a). As discussed in Sec. 2, the planner functions as a critical intermediary that must bridge the conceptual gap between human-readable requirements and coder-executable logic. Specifically, the planner performs two critical transformations: first, it decomposes user inputs into a comprehensive requirement set; then, it designs the coding steps to fulfill these requirements (Dong et al., 2024; Hong et al., 2024):
(3) $\mathcal{R}$ $=\mathcal{A}_{p}^{req}(r)$
(4) $\mathcal{L}$ $=\mathcal{A}_{p}^{logic}(\mathcal{R})$
where $r$ represents the input natural language requirement, $\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ denotes the decomposed requirement set, and $\mathcal{L}=\{l_{1},l_{2},...,l_{m}\}$ represents the coding logic steps expressed in natural language. The generated plan includes both decomposed requirements and coding logic steps:
(5) $p=\{\mathcal{R},\mathcal{L}\}$
Instructed by the coding steps, the coder then translates each step $l_{i}$ into corresponding code blocks:
(6) $c=\mathcal{A}_{c}(\mathcal{L})=\bigcup_{i=1}^{m}\mathcal{A}_{c}(l_{i})$
where $c$ represents the generated code. The planner-coder gap manifests in this transformation, where semantic misalignment occurs between the coding logic and code. Our analysis reveals that this gap accounts for 75.3% of failures (Table 3), highlighting a severe robustness issue.
Fundamentally, the interaction between agents in MASs can be viewed as a cascade of information transformations across agent boundaries, where each agent receives, processes, and transmits information to the next stage. This multi-stage transformation process is inherently susceptible to information loss (Shannon, 1948; Cover and Thomas, 2006). The aforementioned issue of planner-coder gap is further exacerbated by the sequential nature of the transformation process, as there are multiple transformation stages (see Eq. 346). More specifically, each transformation stage can introduce subtle semantic shifts and information loss: the planner might correctly identify what needs to be done but express it in terms that are ambiguous to the coder, or the coder might technically implement the logic steps but miss the underlying intent.
This issue stems from the inherent complexity of inter-agent communication and the structural limitations of current task distribution (Li et al., 2024; Han et al., 2024). Consequently, without robust mechanisms to promote effective communication across agents, MASs remain vulnerable to systematic failures.
Figure 2. Overall workflow of our repairing method. Multi-prompt generation generates various semantically-equivalent versions of user input. The monitor agent conducts plan interpretation and code check to boost communication between planner and coder.
6.2. Overview of Our Repairing Method
To mitigate the information loss discussed in Sec. 6.1, we propose a repairing method that enhances both the diversity of input expressions and the quality of inter-agent communication. Our approach comprises two principal components: multi-prompt generation and monitor agent insertion.
As illustrated in Fig. 2, upon receiving an input requirement, we first mutate the requirement and generate multiple semantic-equivalent expressions to reduce likelihood of MASs misinterpreting specific expressions. For each generation, the planner generates the original plan and forwards it to the monitor, a new agent we introduce to mitigate the planner-coder gap. After receiving the plan from the planner, the monitor agent interprets the plan by providing detailed explanations for core concepts, edge cases, complex logic, relational phrases, and conditional judgments (see Sec. 5), which compensate for the information loss. Subsequently, the coder implements code based on this interpreted plan. In addition, the monitor also checks the implementation to enhance the alignment with the interpreted plan and requests regeneration when misalignments are detected.
6.3. Multi-Prompt Generation
Our analysis in Sec. 4 demonstrated that semantic-preserving mutations can expose robustness issues in MASs. Meanwhile, as the mutation operators explore different expressions of the questions, we observe that these mutations can also clarify ambiguous expressions, potentially improving performance on certain formulations. Leveraging this insight, we propose a multi-generation approach that explores different expressions of the same requirement by reusing the semantic-preserving mutation operators introduced in Table 1.
When the MAS receives an input question, we apply these mutation operators to generate $k$ alternative expressions as generated prompts, with $k+1$ different versions including the original one. During the code generation process, where $n$ represents the total number of generation attempts, we execute the MAS with the original input question and each mutated question for $\frac{n}{k+1}$ times.
This approach enhances robustness by leveraging diverse yet semantically equivalent expressions, allowing the MAS to explore multiple interpretation paths for the same requirement. In our implementation, we employ the MAS’s backend LLM for mutation generation to ensure consistency. We set $k=2$ (three total prompts including the original one) to balance improvement gains with computational efficiency, which also demonstrates generalizability across different MAS architectures and datasets, as shown in Sec. 7.2.
Figure 3. Prompt for the monitor agent. Blue describes the task, brown echoes the five EPs in Sec. 5, while green provides the i/o format and examples.
6.4. Monitor Agent Insertion
Recent studies show that LLMs can provide evaluation highly aligned with humans (Wang et al., 2025; Zheng et al., 2023). Building upon this insight, we introduce a monitor agent between planner and coder to enhance plan quality and strengthen planner-coder alignment. This new agent performs two critical tasks: plan interpretation and code check, serving as a crucial mechanism to compensate for the information loss inherent in the code generation process. Note that this design is motivated by recent findings that specialized agents executing well-defined coordination and evaluation subtasks achieve superior reliability (Fourney et al., 2024; Valmeekam et al., 2023). By focusing on these well-defined, atomic tasks, the monitor exhibits greater robustness compared to the multi-stage transformations between planner and coder that trigger the planner-coder gap.
Plan Interpretation. Despite remarkable efforts in previous work (Dong et al., 2024; Zhang et al., 2024a; Islam et al., 2024) to devise sophisticated prompting structures for planning agents, they primarily focus on choosing the right algorithm without considering the information transmission loss in MASs (see Eq. 34). As discussed in Sec. 6.1, the multi-stage transformation process of MAS introduces information loss, where the planner generates generally correct logic but discards implementation-critical details, leading to incorrect code implementations by the coder.
To effectively address this issue and enhance plan quality, we prompt the inserted monitor agent to interpret the plan generated by the planner to make it more comprehensible to the coder. The monitor agent aims to enhance the clarity and completeness of the planned coding logic and mitigate the planner-coder gap, rather than correcting the semantic content of the plan. Specifically, the monitor interprets the decomposed requirement $\mathcal{R}$ and coding logic $\mathcal{L}$ with detailed explanations and clarifications targeting the five error patterns identified in Sec. 5, producing an interpreted plan $p^{\prime}$ that bridges the semantic gap between planning and coding. Fig. 3 presents an illustrative example of our prompt for the monitor agent. By interpreting the plan $p$ into $p^{\prime}$, the monitor facilitates the restoration of critical implementation details lost in Eq. 34 before code generation, which effectively reduces the information loss between the coding logic and final implementation. We employ few-shot prompting with carefully crafted examples to guide the interpretation process, focusing on the fundamental components of programming plans—such as core concepts, logic flows, and edge cases—rather than being tailored to any specific problem domain. This task-agnostic design provides general applicability across different code generation scenarios.
Clarification. Note that one might consider incorporating the monitor’s interpretation logic directly into the planner’s prompt. However, our proposed approach of a separate monitor agent offers advantages in terms of modularity and scalability. Rather than modifying the planner’s core functionality, this separation of concerns allows the monitor to be seamlessly integrated as a “quality gate” into various existing MAS architectures without altering their fundamental agent logic or task distribution. This design promotes reusability, easier maintenance, and addresses robustness issues across different MAS implementations without requiring architecture-specific modifications. We also follow the practice of consistent backend in a MAS to build the monitor agent with the backend LLMs of the MAS.
Code Check. Although the interpreted plan could preserve crucial details, the information loss between the interpreted plan $p^{\prime}$ and generated code $c$ remains undetected (see Eq. 6), posing a potential issue of misalignment. Therefore, we further incorporate code check as another monitoring task to check semantic alignment between the generated code and the interpreted plan.
Instead of using the testcase-based testing approaches that require dynamic execution with proper running environment, we directly prompt LLMs to perform static inspection and evaluate code compliance with the detailed specifications in $p^{\prime}$, which has been shown to be effective with the rich contextual information provided by the interpreted plan (Tong and Zhang, 2024; Zhuo, 2024). After the coder produces the solution code, we feed both the interpreted plan and the code back to the monitor to check whether the code complies with the plan regarding the different aspects of interpretation, as shown in Fig. 3(c). If the monitor identifies mismatches between code implementation and the plan, it returns the implementation to the coding agent for revision. By checking alignment between the interpreted plan $p^{\prime}$ and generated code $c$, the monitor creates a validation loop that detects and corrects information drift, effectively reducing the information loss that would otherwise accumulate through unchecked transformations. For efficiency, we use zero-shot prompting for code check and limit the process to execute only once.
7. Evaluation
We evaluate our repairing method by answering the following three research questions:
RQ1::
Can our repairing method effectively mitigate failures identified in fuzzing?
RQ2::
How do different components contribute to the overall performance?
RQ3::
Can our repairing method enhance the robustness of MASs against fuzzing?
Specifically, RQ1 first gauges the effectiveness of our method in repairing the failures that are identified in fuzzing. Then, we conduct an ablation study to evaluate the individual contributions of different components within our repairing method in RQ2. In RQ3, we re-run the fuzzing process to evaluate whether our method enhances MASs’ robustness against fuzzing.
7.1. RQ1: Can our repairing method effectively mitigate failures identified in fuzzing?
Table 4. RQ1: Repair results on failures identified through fuzzing.
MAS Backend LLM HumanEval MBPP CodeContest CoderEval
Total Solved Rate Total Solved Rate Total Solved Rate Total Solved Rate
SCCG GPT-3.5 39 30 76.9% 60 48 80.0% 12 9 75.0% 11 9 81.8%
GPT-4o 22 11 50.0% 49 21 42.9% 21 13 61.9% 12 5 41.7%
Deepseek 28 15 53.6% 63 34 53.9% 12 8 66.7% 15 7 46.7%
MetaGPT GPT-3.5 49 37 75.5% 38 23 60.5% 9 5 55.6% 10 8 80.0%
GPT-4o 24 13 54.2% 37 16 43.2% 20 13 65.0% 9 5 55.6%
Deepseek 24 17 70.8% 38 20 53.3% 21 14 66.7% 9 8 88.9%
PairCoder GPT-3.5 15 8 53.3% 25 14 56.0% 22 9 40.9% 19 10 52.6%
GPT-4o 14 7 50.0% 25 16 64.0% 16 7 43.8% 20 14 70.0%
Deepseek 14 7 50.0% 26 14 53.8% 20 8 40.0% 12 9 75.0%
To address RQ1, we evaluate the performance of our repairing method on the failed questions identified during fuzzing. We collect the failures for each MAS with different backend LLMs to form our evaluation dataset. Since all MASs initially failed to solve these questions, the baseline success rate is zero. After applying our repairing method, we re-evaluate the MASs on these questions.
Table 4 presents the compelling repair results: our method enables MASs to successfully solve 40.0%–88.9% of the questions they previously failed. This significant improvement demonstrates that enhanced expression diversity and better planner-coder communication effectively reduce information loss during agent communication and enhance the robustness of MASs. Taking HumanEval-18 as an example, the original plan of MetaGPT (Deepseek) only mentions “please handle the edge case when the substring is empty,” leading to the coder’s misunderstanding. In contrast, the interpretation generated by the monitor further explains the logic of this case by providing examples and a detailed logic flow. Consequently, the coder successfully implements the correct code.
Table 5. Repair ratios for different categories of failures from SCCG (GPT-3.5).
Category EP-1 EP-2 EP-3 EP-4 EP-5 Overall Plan Logic Errors Invalid
Solved (%) 80.0 83.3 80.0 100.0 85.7 83.9 57.1 0.0
The improvement contributed by our method varies across different MAS architectures. As shown in Table 4, SCCG and MetaGPT show exceptional improvement, achieving repair rates exceeding 80%. PairCoder shows more modest gains of 40%–75%. Given that PairCoder incorporates plan selection strategies, the readability and quality of plans generated by PairCoder surpass those of other MASs, therefore suffering less from the planner-coder gap. Similarly, the choice of backend LLM influences repair effectiveness. MASs using GPT-3.5 or Deepseek as backend LLM benefit more from our method compared to those using GPT-4o. These results indicate that our method is more beneficial for less capable MASs with less effective backend LLMs, which typically struggle more with plan comprehension and are more susceptible to the planner-coder gap.
To further understand the strengths of our method, Table 5 presents the repair results for different EPs identified in SCCG (GPT-3.5). Our repairing method demonstrates remarkable effectiveness in handling failures originating from the planner-coder gap, successfully resolving 83.9% of such cases. Notably, the consistently high repair ratios across all five EPs (80.0%–100.0%) indicate that our repairing method serves as a generalizable strategy for bridging the planner-coder gap, regardless of the specific type of error pattern involved. Surprisingly, our method also demonstrates considerable effectiveness on plan logic errors, successfully repairing 57.1% of such cases. Although our approach primarily targets planner-coder gap, this result suggests that the interpretation process may also compensate for the planner’s misunderstandings, as explicit interpretation may inadvertently correct or clarify flawed reasoning in the original plan.
After adopting our method, MASs successfully solve 40.0%–88.9% of failures identified during fuzzing. Our repairing method demonstrates remarkable effectiveness in handling failures originating from the planner-coder gap, successfully repairing the five EPs identified in Sec. 5.
7.2. RQ2: How do different components contribute to the overall performance?
Figure 4. Comparison of repairing performance when removing different components.
Figure 5. Comparison of failures of the original and repaired MASs found in fuzzing.
We conduct ablation studies to investigate the effectiveness of multi-prompt generation and monitor agent insertion in our repairing method. For monitor agent insertion, we separately examine “Plan Interpretation” and “Code Check” to provide in-depth analysis. We present results with GPT-3.5 as the backend LLM, as it represents a typical configuration with other LLMs exhibiting similar patterns.
Fig. 4 shows that removing any component degrades repair performance across all three MASs, indicating that all components contribute meaningfully and our design outperforms all alternative configurations. While disabling “Code Check” causes the smallest performance decrease, enabling “Code Check” still contributes to the overall repair performance and “Only Code Check” contributes approximately 40% of our method’s full repair capability, demonstrating that post-generation validation effectively catches misalignments even without enhanced plan interpretation.
Moreover, our results reveal distinct patterns across MAS architectures, where the relative importance of “Plan Interpretation” versus “Multi-prompt” varies by MAS implementations. In MetaGPT, “Only Plan Interpretation” substantially outperforms “Only Multi-prompt” because MetaGPT generates abstract plans that challenge coder comprehension, where interpretation bridges critical understanding gaps. Conversely, PairCoder benefits more from “Only Multi-prompt” since its plans have better readability, making the diversity of input expressions from multi-prompt generation more beneficial. These findings demonstrate that both strategies are valuable as different MASs exhibit unique flaws: MASs like MetaGPT with abstract plans benefit most from interpretation, while those like PairCoder with clear but rigidly expressed plans gains more from prompt diversity.
Our repairing method achieves optimal performance among all ablation configurations, with each component mitigating different information loss and enhancing the overall repair capability.
7.3. RQ3: Can our repairing method enhance the robustness of MASs against fuzzing?
While RQ1 demonstrates that our repairing method can effectively fix failures identified during the initial fuzzing process, it remains unclear whether our repairing method can prevent or reduce new failures when subjected to continued fuzzing. Therefore, in this RQ, we apply our repairing method to all MASs and re-execute the fuzzing process proposed in Sec. 4. Fig. 5 shows the number of failures discovered from the original MASs and the repaired ones through fuzzing.
From Fig. 5, we observe that on all three MASs, the speed of discovering failures (i.e., the slope of the blue line and the red line) slows down, and identified failures decrease after applying our repairing method, showing that the repaired MASs exhibit superior robustness compared to their original counterparts. Notably, PairCoder (GPT-4o) exhibits the most remarkable robustness improvement against fuzzing, with 85.7% of failures in HumanEval eliminated. SCCG shows stable results across different backend LLMs and datasets, with 28.6%–53.8% decrease in failure numbers. These results indicate that our repairing method can effectively enhance the robustness of MASs against semantic-preserving mutations, resulting in fewer failures discovered during fuzzing.
Across different datasets, MASs demonstrate the most significant robustness improvement on HumanEval, with failures decreasing by 31.2%–85.7%, surpassing other datasets. This superior performance can be attributed to HumanEval’s well-defined questions with clear logic structures. Therefore, our repairing method effectively enhances the coder’s comprehension of the tasks by accurately capturing the question’s intent, thereby substantially improving MAS robustness.
Our repairing method enhances MAS robustness against fuzzing, with up to 85.7% failure number reductions. Repaired MASs exhibit superior robustness on the HumanEval dataset.
8. Discussion
Time Cost. Time efficiency is a critical factor in MASs, as computational overhead directly impacts system scalability and practical deployment (Wu et al., 2023). To understand the time cost of our repairing method, we measure the average execution time per attempt with and without our repairing method. Table 6 presents the experiment results on representative setting (HumanEval, GPT-3.5), as other settings (e.g., datasets, LLMs) show similar trends.
Table 6. Comparison of time cost on HumanEval with GPT 3.5 as backend LLM.
Our method w/o Multi-prompt w/o Monitor Original
SCCG 13.6s 12.9s 11.3s 10.9s
MetaGPT 18.4s 17.6s 15.3s 14.8s
PairCoder 20.1s 19.0s 17.3s 16.4s
In general, the additional time overhead primarily comes from the two components of our repairing method: multi-prompt generation and monitor agent insertion. For multi-prompt generation, since we maintain the same total number of generation attempts $n$, the primary overhead comes only from generating $k$ mutated prompts, which incurs little additional overhead for API calls compared to the main generation process. Besides, the monitor agent introduces two additional API calls per generation attempt, taking merely 2.7s–3.7s across different MASs. We clarify that this overhead is modest in real-world application scenarios, given that a full MAS execution typically takes more than 30 seconds. On the other hand, the adoption of the monitor agent brings substantial robustness improvements, as shown in Sec. 7, making the cost-benefit trade-off favorable.
Threats to Validity. We identified and mitigated several potential threats to validity in our evaluation. The first threat is the representativeness of our experimental setup. To address this, we evaluated three popular MASs (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2024) with three distinct LLMs (OpenAI, 2022; Achiam et al., 2023; Zhu et al., 2024) from different model families across four diverse datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a). The second threat is the reliability of semantic-preserving mutation. To mitigate this threat, we conducted manual verification confirming that 99.2% of mutants preserve semantics. Additionally, we measure the semantic similarity using Sentence-Transformers (Reimers and Gurevych, 2019), whose results show 98.4% average similarity. Thus, the robustness issues revealed by our fuzzing are meaningful. The last threat is subjectivity in manual failure analysis. To ensure reliable implications from failure cases, two software developers with over five years of programming experience independently categorize failure reasons, cross-check the annotation, and resolve all disagreements.
9. Related Work
MASs for Code Generation.  Researchers have designed various MASs with diverse architectures by defining different roles among multiple LLM agents (Hong et al., 2024; Dong et al., 2024; Zhang et al., 2024a; Islam et al., 2024, 2025; Jiang et al., 2024b). For instance, MetaGPT (Hong et al., 2024) employs five distinct agents to simulate a software company’s development workflow, while the Self-Collaboration (Dong et al., 2024) uses three agents for planning, coding, and testing. To enhance plan correctness, Zhang et al. (Zhang et al., 2024a) use clustering to select the optimal plan, and Islam et al. (Islam et al., 2024) employ dynamic traversal to assign confidence scores to plans. However, these methods focus primarily on logical correctness while neglecting the communication gap between planner and coder, which is first identified in this paper. In this work, we propose a repairing method to address the robustness issue of MASs by mitigating the gap.
Testing Code Generation Models. Numerous studies have applied testing methodologies to evaluate code generation LLMs. Wang et al. (Wang et al., 2023a) introduce ReCode, a benchmark for robustness evaluation through systematic perturbation testing, while Mastropaolo et al. (Mastropaolo et al., 2023) conduct empirical studies on GitHub Copilot’s robustness in real-world scenarios. Recent advances have proposed diverse methods to evaluate code generation models under different scenarios (Chen et al., ; Yang et al., 2025; Zhong and Wang, 2024; Chen et al., 2023), like adversarial training (Zhang et al., 2024b) and instruction concretization (Yan et al., 2025). However, none of these works focus on MASs and their robustness remains under-explored. We conducted the testing-based study over MAS robustness and identified the planner-coder gap as the main cause of robustness issues.
10. Conclusion
This paper presents the first comprehensive study on the robustness of MASs for code generation, revealing their robustness issue on semantically equivalent mutations. We identify the “planner-coder gap” as the root cause and propose a novel repair method incorporating multi-prompt generation and monitor agent insertion. Our findings provide essential insights for developing more reliable MASs and establish a foundation for future robustness research in this domain.